When we say a TV is 42 in., we mean the diagonal is 42 in. For a 42-in TV, the width is 15 in. more than the height. Find the dimensions of a 42-in. TV.
Answer by rwm(914)

42^2=x^2+(x+15)^2
1764 = x^2+30x+225
x=21.2359
x+15=36.2359